OpenStudy (anonymous):

A boy is 24 years younger than his father. In 2 years time the sum of their ages will be 40. Find the present ages of father and son.

OpenStudy (anonymous):

we have two equations and 2 unknowns. we can solve for both you just need the two equations and re arrange them accoridingly

OpenStudy (anonymous):

2(f+b)=40**

OpenStudy (anonymous):

And thank you (:

OpenStudy (anonymous):

yeah sorry. If you caught my error you should be good

OpenStudy (anonymous):

Genuine, I disagree with that second equation.

OpenStudy (anonymous):

The sum of their ages 2 years from now is not 2(f+b)

OpenStudy (anonymous):

hmm let me check

OpenStudy (anonymous):

wait, yes @SmoothMath you're right

OpenStudy (anonymous):

yeah it's phrased weird in two years time the sum of their ages.

OpenStudy (anonymous):

I mean, just think about a specific case. If I'm 10 and you're 12, the sum of our ages two years from now will not be 44.

OpenStudy (anonymous):

how would you write that out instead It's still not clear.

OpenStudy (anonymous):

I don't know if im right but using Romero's equations I worked out that father was 22 years old?

OpenStudy (anonymous):

Father's age 2 years from now: f+2 Boy's age 2 years from now: b+2 Sum of those ages: (f+2) + (b+2) = f+ b +4

OpenStudy (anonymous):

I read it wrong. It's not using the word "time" to represent multiplication

OpenStudy (anonymous):

Oooh. I see how you could misread that.

OpenStudy (anonymous):

Easy mistake. But yeah, the actual system of equations we get is: f-b = 24 f+b+4 = 40

OpenStudy (anonymous):

Oh, i've got it now (: Thank you guys